OpenStudy (anonymous):

Use less than, equal to, or greater than to complete the statement. The measure of each exterior angle of a regular 7-gon is ____ the measure of each exterior angle of a regular 5-gon. cannot tell equal to less than greater than

OpenStudy (saifoo.khan):

let me solve.

OpenStudy (anonymous):

ok do it :)

OpenStudy (saifoo.khan):

formulae is 360/n

OpenStudy (anonymous):

ok

OpenStudy (saifoo.khan):

7sided = 360/7 = 51.4 5 sided = 360/n = 72

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

so it would be D?

OpenStudy (saifoo.khan):

C

OpenStudy (anonymous):

tankyo!!

OpenStudy (saifoo.khan):

uw